# Concept – An Evaluation Protocol on Conversation Recommender Systems with System- and User-centric Factors

### v1

@@ Abstract
Concept – An Evaluation Protocol on Conversation Recommender Systems with System- and User-centric Factors
[…]
@@ 1 Introduction
By applying Concept, we can evaluate and analyze the strengths, weaknesses, and potential risks of off-the-shelf CRS models [^1] We conduct evaluations using both humans and LLM, and found highly correlated results of the two.. A total of 6720 conversation data is recorded to collect the interactions between off-the-shelf CRS and simulated users who demonstrate different personas and preferences. Existing works agree on the insurmountable performance achieved by ChatGPT-based CRS. However, current CRS models, even enhanced by "omnipotent" ChatGPT, still encounter numerous challenges, falling short of practical usability in particular. 1) They struggle to express genuine responses without hallucination or deceit, often introducing non-existent items into conversations and present to users. 2) They lack of self-awareness of its identity, facing difficulties in producing both persuasive and honest explanations, which is prominent in the ChatGPT-based CRS model, where explanations are highly convincing but frequently contain illusory details, misleading users in believing that these items align with their preferences. 3) They encounter issues in offering reliable recommendations as they are sensitive to contextual nuances. Even slight alterations in user wording may result in entirely different recommendations. 4) They lack proficiency in catering to diverse users without prior coordination, failing to dynamically adjust their behavior to align with each user’s distinct personas. It is observed that the ChatGPT-based CRS model tends to employ deceptive tactics to persuade optimistic users to accept recommendations. This underscores the importance of aligning CRS with human values and advocating its ethical use.
With Concept, we provide an overview for researchers as a reference guidance for evaluating CRS and laying the foundation for CRS enhancement. The main contributions as follows:
@@ 2 Related Work
Our research focuses on the CRS evaluation, with LLM-based methods in particular. Hence, our literature review discusses the same scope of research. Most relevant works are summarized below, and more detailed review is attached in Appendix B.
@@ 3.1 Factors Categorization
Interdisciplinary research on conversational AI Chaves and Gerosa (2021); Reeves and Nass (1996); Fogg (2003) groups its characteristics that impact the user experience into three types. Concept then customizes them for CRS in this section, as depicted in Figure 1. For details on the customization, refer to Appendix C.
3.1.1 Recommendation Intelligence
@@
From a system-centric factors, this requires CRS to learn from conversations and evolve toward making recommendations as the conversation advances  Chen et al. (2019); Ma et al. (2020); Zhou et al. (2021). In this regard, our categorization encompasses two primary abilities:
  • Quality. CRS should provide precise recommendations using minimal conversation turns, which are the crucial aspects that influence user satisfaction Siro et al. (2023); Gao et al. (2021). We emphasize the importance of the user acceptance rate, reflecting the practical effectiveness of the recommendations.
@@
    In our experiments, we utilize computational metrics for automatic evaluation, as seen in prior studies Wang et al. (2023d); Zhang et al. (2023); Yu et al. (2023). These metrics encompass $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate, and average turns ($AT$) needed to achieve successful recommendations.
  • Reliability. CRS should deliver robust and consistent recommendations that account for contextual nuances. In practical situations, users with similar preferences may express themselves differently. It would diminish the user experience and pose a disruption for critical applications if inconsistent items were recommended for two similar user responses Tran et al. (2021); Oh et al. (2022). In less critical cases, if two recommended items are inconsistent but align with user preferences, they are often viewed as diverse recommendations, creating opportunities for relevant but less popular items Yang et al. (2021).
@@
    To evaluate reliability, we generate sets of user response pairs with similar meanings using ChatGPT paraphrasing. Given a user response pair [$u_{1}$, $u_{2}$], we define four computational metrics: the rate of Consistent Action, indicating whether the CRS constantly provides recommendations based on $u_{1}$ and $u_{2}$; the rate of Consistent Recommendation, which assesses if CRS recommends the same items given two user responses $u_{1}$ and $u_{2}$; and the rates of Diversity, which evaluate whether the recommended items, even if inconsistent, align with user preferences. We refer to it as Sensitivity when CRS provides inconsistent and inaccurate recommendations that do not align with user preferences.
3.1.2 Social Intelligence
@@
It requires CRS to produce adequate social behavior for the recommendation during the conversations. As evidenced by the Media Equation Theory [^3] Users tend to engage with the machine in a manner that mirrors person-to-person conversations. Reeves and Nass (1996); Fogg (2003), users have high expectations for CRS to act cooperatively and be aware of user’s social needs during the conversation, facilitating the design of CRS with perceived humanness Jacquet et al. (2018, 2019). Our categorization encompasses two abilities.
@@ 4. Relevance
Table 1: Summary of the evaluation taxonomy, descriptions of abilities, and evaluation metrics in Concept. The LLM-based evaluator is used for ability-specific scoring, whereas computational metrics are used for the rest (cf. Table 18 in Appendix for details). We adjust the score to a scale of 1 to 5 when needed.
3.1.3 Personification
Personification requires CRS to perceive the identity of itself and the personality representation of users. This involves CRS being self-aware of its role as a conversational recommendation system tailored for the general public.
  • Identity. CRS should be self-aware of its identity and operate within its designated scope, differentiating itself from sales systems or other types. This ensures that CRS effectively carries out its identity by offering persuasive explanations to boost user acceptance Jannach et al. (2021); Zhou et al. (2022), avoiding resorting to misleading strategies (i.e., sales pitches with deceptive tactics). Otherwise, user trust and loyalty are hindered Gkika and Lekakos (2014). Using misleading strategies also violates the maxim of sincerity (cf. Section 3.1.2). We emphasize the need to integrate self-awareness into CRS to mitigate any deceptive behavior.
    We evaluate the identity by calculating the proportion of non-deceptive explanations. We also resort to the LLM-based evaluator to score the persuasiveness of recommendation explanations.
  • Coordination. CRS should be proficient in serving users possessing various personas without prior coordination. Attributing personality to a conversational AI ensures that its behaviors stand in agreement with the users’ expectations in a particular context Chaves and Gerosa (2021). This becomes particularly challenging for CRS, as it frequently serves users with diverse personas in real-world scenarios Thompson et al. (2004). As a result, CRS needs to exhibit various personalities and adapt its behavior to suit different users Katayama et al. (2019); Svikhnushina et al. (2021); Zhou et al. (2020). A crucial ability is to proficiently serve users without prior coordination.
    We create simulations of users with different personas and evaluate the performance of the CRS across all the abilities mentioned. To assess coordination, we initially calculate the range [^4] The range is more effective than the standard deviation in highlighting the variability of the CRS across different users. and mean of the CRS’s scores for each specific ability across different users. Then, we divide the range by the mean to determine the coordination score of the CRS for that specific ability. The overall coordination score is subsequently calculated as the average across all abilities.
@@ 4.1 Experimental Setup
CRS Model. Following (Wang et al., 2023d), we present a comparative evaluation and analysis towards representative and SOTA CRS models, including KBRD Chen et al. (2019), BARCOR Wang et al. (2022a), UNICRS Wang et al. (2022b), and CHATCRS Wang et al. (2023d). Refer to Appendix A for implementation details.
Figure 2: An overview of the evaluation of off-the-shelf CRS models. The practical usability of those models are still hindered.
4.2 Evaluation Using Concept
Following the recent study on CRS Wang et al. (2023d), Concept resorts to an LLM-based user simulator and evaluator for cost-effective evaluation, together with computational metrics. We summary the taxonomy and evaluation metrics in Table 1. See Appendix A.2 for evaluation details.
LLM-based User Simulator. Our user simulator has unique personas and preferences. The personas are generated by prompting ChatGPT in a zero-shot manner, following (Wang et al., 2023a), while the preferences are defined using attributes from two benchmark datasets (refer to Section 4.1). In addition, Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer (2023). This is achieved by prompting the simulator to first assess its current mental state before generating responses, eliciting reflection on its predefined personality traits and social interactions. See Appendix A.2.2 for more implementation details on the simulator.
LLM-based Evaluator. Following previous studies Ye et al. (2023); Wang et al. (2023c), we employ the instance-wise evaluator. Building upon prior research Wang et al. (2023d); Liu et al. (2023d), we prompt the evaluator with fine-grained scoring rubrics to eliminate the scoring bias [^6] See Appendix D for our reliability analysis.. The evaluator assigns a score ranging from 1 to 5 to the conversation data using ability-specific rubrics, each accompanied by a corresponding description. For the generation of fine-grained rubrics, we follow the approach of previous works Saha et al. (2023a); Li et al. (2023b) and employ ChatGPT to produce a set of evaluation criterion, which are then refined by humans. See Appendix A.2.3 for details.
4.3 Evaluating off-the-shelf CRS models
Figure 2 presents an overview of the results across the six primary abilities, averaged across two benchmark datasets [^7] For results on each benchmark, refer to Appendix E.. These results suggest that CHATCRS has made significant progress in cooperation, social awareness, and recommendation quality while losing advantages in identity. We will continue the current discussion in subsequent sections and provide further analysis.
4.3.1 Recommendation-centric Evaluation
@@
CHATCRS stands out as the leading CRS model, delivering higher-quality recommendations. As indicated in Table 2, most CRS models have demonstrated improved success rates (SR) from the conversation module’s perspective, indicating the effectiveness of optimizing conversation and recommendation modules simultaneously. However, we observed lower SR values for BARCOR on the Redial dataset compared to those from the recommendation module perspective, which is linked to the inclusion of non-existent items by its conversation module, such as "The Adventures of Milo and Ours" (should be ’Otis’) and "The Prestigige" (should be ’Prestige’). Despite this, their recall [^8] We only specifies user preferences, leading to a wide range of target items for users, resulting in a low recall value. and success rate (SR) values still fall short. In addition, the user acceptance rate of CRS models other than CHATCRS is notably low, possibly due to users’ intolerance towards poor recommendations and the absence of persuasive recommendation explanations, which we will investigate on in further studies. Instead, CHATCRS stands out as the leading CRS model, with significant performance improvements, conforming to recent findings Wang et al. (2023d). This may be attributed to two factors: 1) The text-embedding-ada-002 is much stronger in terms of translating conversation context and user preferences into embeddings, which helps enhance the recommendation module significantly. 2) CHATCRS are persuasive in convincing users to accept the recommendations. However, such a high user acceptance rate is largely based on deceptive tactics, a topic we will delve into later.
Figure 3: Reliability evaluation on CHATCRS, which is sensitive to contextual nuance.
@@
Current CRS strives to offer reliable recommendations due to its sensitivity to contextual nuance. We gauge the model reliability by evaluating how off-the-shelf CRS models handle user responses that are semantically similar but expressed differently. Results of CHATCRS [^9] More results are in Table 12. are presented in Figure 3, which show promising action consistency of CHATCRS. It can make consistent recommendations when presented with two semantically similar user responses, achieving a consistency rate of over 99%. However, the average recommendation consistency rate is much lower, at only 51.58%, which means that 48.42% of the time, if the user alters their wording slightly, CHATCRS will recommend entirely different items. To further explore its impact, we conducted a thorough analysis of the relationship between the attributes of the recommended items and user preferences. We discovered that only about 12%-17% of the recommended items align with user preferences, despite being inconsistent. This may increase chances for relevant but less popular items. As a consequence, the majority of recommendation results do not align with user preferences, demonstrating the sensitivity of current CRS to contextual nuances and the negative impact on user experience.
@@ 4.3.2 Social-centric Evaluation
Figure 5: Coordination evaluation on users with various personas (marked with dashed boxes). CHATCRS is hard to cater to diverse users. Others consistently show low performance levels that don’t vary much across different users.
@@ 4.3.3 Personification-centric Evaluation
Figure 6: Identity and sincerity evaluation. CHATCRS offers persuasive yet dishonest explanations.
@@
Reliability of Concept. We have established fixed values for the temperature and seed parameters to guarantee replicability of our LLM-based simulator and evaluator. We show that Concept scoring is unbiased and aligns with the results of human evaluations (cf. Appendix D).

### v2

# Concept – An Evaluation Protocol on Conversation Recommender Systems with System-centric and User-centric Factors

@@ 1 Introduction
By applying Concept, we can evaluate and analyze the strengths, weaknesses, and potential risks of off-the-shelf CRS models [^1] We conduct evaluations using both humans and LLM, and found highly correlated results of the two.. A total of 6720 conversation data is recorded to collect the interactions between off-the-shelf CRS and simulated users who demonstrate different personas and preferences. Existing works agree on the insurmountable performance achieved by ChatGPT-based CRS. However, current CRS models, even enhanced by "omnipotent" ChatGPT, still encounter numerous challenges, falling short of practical usability in particular. 1) They struggle to express genuine responses without hallucination or deceit, often introducing non-existent items into conversations and present to users. 2) They lack of self-awareness of its identity, facing difficulties in producing both persuasive and honest explanations, which is prominent in the ChatGPT-based CRS model, where explanations are highly convincing but frequently contain illusory details, misleading users in believing that these items align with their preferences. 3) They encounter issues in offering reliable recommendations as they are sensitive to contextual nuances. Even slight alterations in user wording may result in entirely different recommendations. 4) They lack proficiency in catering to diverse users without prior coordination, failing to dynamically adjust their behavior to align with each user’s distinct personas. It is observed that the ChatGPT-based CRS model tends to employ deceptive tactics to persuade optimistic users to accept recommendations. This underscores the importance of aligning CRS with human values and advocating its ethical use. With Concept, we provide an overview for researchers as a reference guidance for evaluating CRS and laying the foundation for CRS enhancement. The main contributions as follows:
@@ 2 Related Work
Our research focuses on the CRS evaluation, with LLM-based methods in particular. Hence, our literature review discusses the same scope of research. Most relevant works are summarized below, and more detailed review is attached in Appendix C.
@@ 3.1 Factors Categorization
Interdisciplinary research on conversational AI Chaves and Gerosa (2021); Reeves and Nass (1996); Fogg (2003) groups its characteristics that impact the user experience into three types. Concept then customizes them for CRS in this section, as depicted in Figure 1.
@@
Factor 1: Recommendation Intelligence. From a system-centric factors, this requires CRS to learn from conversations and evolve toward making recommendations as the conversation advances  Chen et al. (2019); Ma et al. (2020); Zhou et al. (2021). In this regard, our categorization encompasses two primary abilities:
@@
  • Quality. CRS should provide precise recommendations using minimal conversation turns, which are the crucial aspects that influence user satisfaction Siro et al. (2023); Gao et al. (2021). We emphasize the importance of the user acceptance rate, reflecting the practical effectiveness of the recommendations. In our experiments, we utilize computational metrics for automatic evaluation, as seen in prior studies Wang et al. (2023d); Zhang et al. (2023); Yu et al. (2023). These metrics encompass $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate, and average turns ($AT$) needed to achieve successful recommendations.
@@
  • Reliability. CRS should deliver robust and consistent recommendations that account for contextual nuances. In practical situations, users with similar preferences may express themselves differently. It would diminish the user experience and pose a disruption for critical applications if inconsistent items were recommended for two similar user responses Tran et al. (2021); Oh et al. (2022). In less critical cases, if two recommended items are inconsistent but align with user preferences, they are often viewed as diverse recommendations, creating opportunities for relevant but less popular items Yang et al. (2021). To evaluate reliability, we generate sets of user response pairs with similar meanings using ChatGPT paraphrasing. Given a user response pair [$u_{1}$, $u_{2}$], we define four computational metrics: the rate of Consistent Action, indicating whether the CRS constantly provides recommendations based on $u_{1}$ and $u_{2}$; the rate of Consistent Recommendation, which assesses if CRS recommends the same items given two user responses $u_{1}$ and $u_{2}$; and the rates of Diversity, which evaluate whether the recommended items, even if inconsistent, align with user preferences. We refer to it as Sensitivity when CRS provides inconsistent and inaccurate recommendations that do not align with user preferences.
@@
Factor 2: Social Intelligence. It requires CRS to produce adequate social behavior for the recommendation during the conversations. As evidenced by the Media Equation Theory [^3] Users tend to engage with the machine in a manner that mirrors person-to-person conversations. Reeves and Nass (1996); Fogg (2003), users have high expectations for CRS to act cooperatively and be aware of user’s social needs during the conversation, facilitating the design of CRS with perceived humanness Jacquet et al. (2018, 2019). Our categorization encompasses two abilities.
@@ 4. Relevance
Table 1: Summary of the evaluation taxonomy, descriptions of abilities, and evaluation metrics in Concept. LLM-based evaluator is used for ability-specific scoring, whereas computational metrics are used for the rest (cf. Table 18 for details). We adjust the score to a scale of 1 to 5 when needed.
Factor 3: Personification. Personification requires CRS to perceive the identity of itself and the personality representation of users. This involves CRS being self-aware of its role as a conversational recommendation system tailored for the general public.
  • Identity. CRS should be self-aware of its identity and operate within its designated scope, differentiating itself from sales systems or other types. This ensures that CRS effectively carries out its identity by offering persuasive explanations to boost user acceptance Jannach et al. (2021); Zhou et al. (2022), avoiding resorting to misleading strategies (i.e., sales pitches with deceptive tactics). Otherwise, user trust and loyalty are hindered Gkika and Lekakos (2014). Using misleading strategies also violates the maxim of sincerity (cf. Section 3.1). We emphasize the need to integrate self-awareness into CRS to mitigate any deceptive behavior. We evaluate the identity by calculating the proportion of non-deceptive explanations. We also resort to the LLM-based evaluator to score the persuasiveness of recommendation explanations.
  • Coordination. CRS should be proficient in serving users possessing various personas without prior coordination. Attributing personality to a conversational AI ensures that its behaviors stand in agreement with the users’ expectations in a particular context Chaves and Gerosa (2021). This becomes particularly challenging for CRS, as it frequently serves users with diverse personas in real-world scenarios Thompson et al. (2004). As a result, CRS needs to exhibit various personalities and adapt its behavior to suit different users Katayama et al. (2019); Svikhnushina et al. (2021); Zhou et al. (2020). A crucial ability is to proficiently serve users without prior coordination. We create simulations of users with different personas and evaluate the performance of the CRS across all the abilities mentioned. To assess coordination, we initially calculate the range [^4] The range is more effective than the standard deviation in highlighting the variability of the CRS across different users. and mean of the CRS’s scores for each specific ability across different users. Then, we divide the range by the mean to determine the coordination score of the CRS for that specific ability. The overall coordination score is subsequently calculated as the average across all abilities.
@@ 4.1 Experimental Setup
CRS Model. Following (Wang et al., 2023d), we present a comparative evaluation and analysis towards representative and SOTA CRS models, including KBRD Chen et al. (2019), BARCOR Wang et al. (2022a), UNICRS Wang et al. (2022b), and CHATCRS Wang et al. (2023d).
4.2 Evaluation Using Concept
Following the recent study on CRS Wang et al. (2023d), Concept resorts to an LLM-based user simulator and evaluator for cost-effective evaluation, together with computational metrics. We summary the taxonomy and evaluation metrics in Table 1. See Appendix B.2 for evaluation details.
LLM-based User Simulator. Our user simulator has unique personas and preferences. The personas are generated by prompting ChatGPT in a zero-shot manner, following (Wang et al., 2023a), while the preferences are defined using attributes from two benchmark datasets (refer to Section 4.1). In addition, Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer (2023). This is achieved by prompting the simulator to first assess its current mental state before generating responses, eliciting reflection on its predefined personality traits and social interactions. See Appendix B.2.2 for more implementation details on the simulator.
LLM-based Evaluator. Following previous studies Ye et al. (2023); Wang et al. (2023c), we employ the instance-wise evaluator. Building upon prior research Wang et al. (2023d); Liu et al. (2023d), we prompt the evaluator with fine-grained scoring rubrics to eliminate the scoring bias. The evaluator assigns a score ranging from 1 to 5 to the conversation data using ability-specific rubrics, each accompanied by a corresponding description. For the generation of fine-grained rubrics, we follow the approach of previous works Saha et al. (2023a); Li et al. (2023b) and employ ChatGPT to produce a set of evaluation criterion, which are then refined by humans. See Appendix B.2.3 for details.
4.3 Evaluating off-the-shelf CRS models
Figure 2: Evaluation overview.
Figure 2 presents an overview of the results across the six primary abilities, averaged across two benchmark datasets [^6] For results on each benchmark, refer to Appendix F.. These results suggest that CHATCRS has made significant progress in cooperation, social awareness, and recommendation quality while losing advantages in identity. We will continue the current discussion in subsequent sections and provide further analysis.
4.3.1 Recommendation-centric Evaluation
@@
CHATCRS stands out as the leading CRS model, delivering higher-quality recommendations. As indicated in Table 2, most CRS models have demonstrated improved success rates (SR) from the conversation module’s perspective, indicating the effectiveness of optimizing conversation and recommendation modules simultaneously. However, we observed lower SR values for BARCOR on the Redial dataset compared to those from the recommendation module perspective, which is linked to the inclusion of non-existent items by its conversation module, such as "The Adventures of Milo and Ours" (should be ’Otis’) and "The Prestigige" (should be ’Prestige’). Despite this, their recall [^7] We only specifies user preferences, leading to lots of target items for users, resulting in a low recall value. and success rate (SR) values still fall short. In addition, the user acceptance rate of CRS models other than CHATCRS is notably low, possibly due to users’ intolerance towards poor recommendations and the absence of persuasive recommendation explanations, which we will investigate on in further studies. Instead, CHATCRS stands out as the leading CRS model, with significant performance improvements, conforming to recent findings Wang et al. (2023d). This may be attributed to two factors: 1) The text-embedding-ada-002 is much stronger in terms of translating conversation context and user preferences into embeddings, which helps enhance the recommendation module significantly. 2) CHATCRS are persuasive in convincing users to accept the recommendations. However, such a high user acceptance rate is largely based on deceptive tactics, a topic we will delve into later.
Figure 3: CHATCRS is sensitive to contextual nuance.
@@
Current CRS strives to offer reliable recommendations due to its sensitivity to contextual nuance. We gauge the model reliability by evaluating how off-the-shelf CRS models handle user responses that are semantically similar but expressed differently. Results of CHATCRS are presented in Figure 3, which show promising action consistency of CHATCRS. It can make consistent recommendations when presented with two semantically similar user responses, achieving a consistency rate of over 99%. However, the average recommendation consistency rate is much lower, at only 51.58%, which means that 48.42% of the time, if the user alters their wording slightly, CHATCRS will recommend entirely different items. To further explore its impact, we conducted a thorough analysis of the relationship between the attributes of the recommended items and user preferences. We discovered that only about 12%-17% of the recommended items align with user preferences, despite being inconsistent. This may increase chances for relevant but less popular items. As a consequence, the majority of recommendation results do not align with user preferences, demonstrating the sensitivity of current CRS to contextual nuances and the negative impact on user experience.
@@ 4.3.3 Personification-centric Evaluation
Figure 5: CHATCRS offers persuasive yet dishonest explanations.
Figure 6: Coordination evaluation on users with various personas (marked with dashed boxes). CHATCRS is hard to cater to diverse users, while others consistently show low performance levels.
@@
Figure 7: Length bias evaluation.
4.4 Reliability of Concept
This section aim to assess the reliability of Concept in terms of its replicability and potential evaluation bias.
Replicability Analysis. We have established fixed values for the temperature and seed parameters to guarantee replicability of our LLM-based simulator and evaluator.
Bias Analysis. The LLM-based evaluation is prone to biases Wang et al. (2023e). To address this, we build upon previous research Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d) and minimize these potential biases by employing detailed scoring rubrics. In this study, we examine the influence of length and self-enhancement biases [^8] When assessing CRS, Concept does not entail making decisions based on a specific group of candidates, thus freeing it from any position bias., and demonstrate that Concept scoring is unbiased and aligns with the results of human evaluations, consistent with earlier discoveries Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d).
  • Length bias [^9] LLMs have a preference for longer responses Wu and Aji (2023).. As shown in Figure 7, we require CHATGPT to provide an overall performance score based on all the ability-specific scores (i.e., Y-axis, Overall Performance Score) and then plot the relationship between various CRS reply lengths (i.e., X-axis, Avg. Response Length) and the scores. Our findings show that Concept remains unaffected by the length bias. Despite CHATCRS tending to produce lengthier responses, this does not imply that longer responses will yield higher scores.
  • Self-enhancement bias [^10] Bias of CHATGPT to favor high scores for its generated content Wang et al. (2023c).. Our human evaluation demonstrates that the Concept evaluation results are consistent and align with human assessments (cf. Appendix E for more details). Specifically, the human evaluation results based on the conversation data and the LLM-based evaluation results are closely related, with a correlation coefficient of 61.24% and Krippendorff’s alpha of 53.10%. This indicates the reliability of our LLM-based evaluation, which is consistent with previous findings Wang et al. (2023d).